OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. x^2 + 4x + y^2 − 6y = −4

OpenStudy (anonymous):

@ParthKohli can you help?

OpenStudy (sidsiddhartha):

first complete the square

OpenStudy (anonymous):

What do you mean complete the square?

OpenStudy (sidsiddhartha):

i mean make it like (x-A)^2+(y-B)^2=C^2

OpenStudy (anonymous):

I'm sorry. I am confused.

OpenStudy (sidsiddhartha):

I'LL look like (x+2)^2+(y-3)^2=(3)^2 so centre=(-2,3) radius=3

OpenStudy (anonymous):

oh okay. im sorry. This kind of math has always confused me.

OpenStudy (sidsiddhartha):

 now u get it?